### v3

# Learning graph geometry and topology using dynamical systems based message-passing

Abstract
In this paper we introduce DYMAG: a message passing paradigm for GNNs built on the expressive power of continuous, multiscale graph-dynamics. Standard discrete-time message passing algorithms implicitly make use of simplistic graph dynamics and aggregation schemes which limit their ability to capture fundamental graph topological properties. By contrast, DYMAG makes use of complex graph dynamics based on the heat and wave equation as well as a more complex equation which admits chaotic solutions. The continuous nature of the dynamics are leveraged to generate multiscale (dynamic-time snapshot) representations which we prove are linked to various graph topological and spectral properties. We demonstrate experimentally that DYMAG achieves superior performance in recovering the generating parameters of Erdös-Renyi and stochastic block model random graphs and the persistent homology of synthetic graphs and citation network. Since the behavior of proteins and biomolecules is sensitive to graph topology and exhibits important structure at multiple scales, we find that DYMAG outperforms other methods at predicting salient features of various biomolecules.
1 Introduction
Graphs provide a flexible way to encode relational, structural, and geometric relations in data, thereby making them a broadly applicable modeling framework in a multitude of application domains. As such, recent years have seen a proliferation in machine learning methods for graph-structured data, mostly notably Graph Neural Networks (GNNs), which have achieved success on numerous tasks such as graph classification, node classification, and link prediction [6, 45]. However, GNNs are limited in their ability to capture graph structure and topology[43]. Concretely, this results in decreased performance on tasks that depend heavily on graph structure (rather than informative node features), such as those arising in molecular and protein property prediction [22]. In this paper, we directly study graph topological property recovery and propose a method for recovering such properties based on a novel form of message passing rooted in continuous, multiscale dynamics.
Examining the interplay between partial differential equations (PDEs) and various graph learning methods provides a unified framework that enables us to identify two limitations of existing graph neural networks which we ameliorate using DYMAG. As discussed in Chamberlain et al. [8], the message passing operation of various graph neural networks can be understood as single-step schemes which make discrete approximations to solutions of a diffusive partial differential equation (PDE). This makes GNNs prone to oversmoothing, limits the viable depth of the network, and leads to the exponential convergence of node features. Hence the implicit choice of diffusion dynamics leads to smooth node features that limit the expressivity of the method, and leads us to our first observation:
(1) A large class of graph learning methods are based around simplistic dynamics.
A general strategy in GNN architectures is to perform message passing for multiple rounds and then aggregate the final node features to generate node or graph level representations. As noted in [8], these message passing operations can be understood as corresponding to the discretization of an underlying diffusive PDE (such as the heat equation) is which is based on the local topology of the graph and produces progressively smoother representations as time increases, leading to an issue commonly referred to as the oversmoothing problem. Moreover, in a standard message passing network, the final depth of message passing (corresponding to the final time step) defines both the reach and the effective time scale of the graph representation. Thus, these methods are limited to creating only a single snapshot of the underlying continuous dynamics and generating a graph representation at one scale of resolution. Hence we identify a second limitation:
(2) During message passing, the implicit sampling of an underlying graph PDE at a single time point generates a representation at a single scale. However, multiple scales of resolution may contain important information about network structure and topology.
We propose Dynamic Multiscale Aggregation on Graphs (DYMAG) to directly address these two issues. To address point (1), DYMAG utilizes complex, continuous-time dynamics. We demonstrate in Section 3.1 that these dynamics are closely linked to several interesting graph topological and spectral properties. This motivates our choice of dynamics, as the solutions (representations) generated from our chosen dynamics display convergent, periodic, and chaotic dynamics which are important for our expressiveness. To address point (2), we sample the underlying continuous PDE at multiple time points to capture the dynamic evolution of node features and provide a multiscale view of the graph at a time-parameterized continuum of scales. The theoretical motivations for DYMAG are given in Section 3.1 and the explicit architecture is described in Section 3.2.
We next experimentally demonstrate the ability of DYMAG to overcome the limitations of standard GNNs described in point (1) and (2) concerning their inability to capture graph topological properties and multiscale views of graph properties. We first validate our method through several experiments on synthetic and citation network datasets described in Section 4.1. We then demonstrate the superior performance of DYMAG on a variety of biomolecular tasks where graph topology and multiscale interactions are of critical importance.
To summarize, our main contributions are as follows:
  1. 1.
    We introduce DYMAG, a novel method for message passing based on continuous, complex graph dynamics which is motivated by our theoretical results linking graph dynamics to graph topological and spectral properties.
  2. 2.
    We utilize the continuous nature of the underlying graph dynamics and extract multiple time point snapshots to generate rich multiscale representations.
  3. 3.
    We show that our method is able to predict geometric and topological properties of the network such as curvature and extended persistence image better than standard message passing networks. We additionally validate its superior performance on recovering dihedral angles and bond lengths in proteins, and the total polar surface area (TPSA) and the number of aromatic rings in molecules.
2 Background: Dynamics
2.1 Heat and Wave Dynamics
For an undirected graph $G=(V,E)$ with vertices $V=\{v_{1},\ldots,v_{n}\}$, we denote $L$ the (normalized or unnormalized) graph Laplacian with eigenvectors $\boldsymbol{\nu}_{1},\ldots,\boldsymbol{\nu}_{n}$ and eigenvalues $0=\lambda_{1}\leq\lambda_{2}\leq\ldots\leq\lambda_{n}$. For a time-varying function $u:V\times[0,\infty)\to\mathbb{R}$, we say that $u(v,t)$ solves the heat equation with a initial data $\mathbf{x}:V\rightarrow\mathbb{R}$ if
$-Lu_{H}=\partial_{t}u_{H},\quad u_{H}(v,0)=\mathbf{x}(v),$ (1)
and say that $u$ solves the wave equation with initial data $\mathbf{x}$ and a initial velocity $\mathbf{y}:V\rightarrow\mathbb{R}$ if
$-L_{v}u_{W}=\partial^{2}_{t}u_{W},\>\>\>u_{W}(v,0)=\mathbf{x}(v),\>\>\>\partial_{t}u_{W}(v,0)=\mathbf{y}(v).$ (2)
If $G$ is connected, solutions to these two equations are given explicitly by
$u_{H}(v,t)$ $=\sum_{i=1}^{n}e^{-t\lambda_{i}}\langle\boldsymbol{\nu_{i}},\mathbf{x}\rangle\boldsymbol{\nu_{i}}(v),\quad\text{and}$ (3)
$u_{W}(v,t)$ $=\sum_{i=1}^{n}\cos(\sqrt{\lambda_{i}}t)\langle\boldsymbol{\nu}_{i},\mathbf{x}\rangle\boldsymbol{\nu}_{i}(v)+t\langle\boldsymbol{\nu}_{1},\mathbf{y}\rangle\boldsymbol{\nu}_{1}+\sum_{i=2}^{n}\frac{1}{\sqrt{\lambda_{i}}}\sin(\sqrt{\lambda_{i}}t)\langle\boldsymbol{\nu}_{i},\mathbf{y}\rangle\boldsymbol{\nu}_{i}(v).$ (4)
Importantly, we note that by imitating the same argument as in Remark 1 of Chew et al. [9], the solutions $u_{H}(v,t)$ and $u_{W}(v,t)$ defined in Eqns. 3 and 4 do not depend on the choice of orthonormal basis for the graph Laplacian. Additionally, the definitions of the heat and wave equations in Eqns. 3 and 4 apply to disconnected graphs. The solution to the heat equation in Eqn. 3 is unchanged in the case where $G$ is disconnected and it is straightforward to modify Eqn. 4 for a disconnected graph.
2.2 Chaotic Dynamics
The linear dynamics considered in the previous section give rise to either periodic solutions, or solutions which converge exponentially fast to an equilibrium state. By contrast, certain nonlinear dynamical systems exhibit chaos–a behavior that may be informally summarized as “aperiodic long-term behavior in a deterministic system that exhibits sensitive dependence on initial conditions" [38]. This sensitivity to initial condition is captured by the Lyapunov exponent. For two separate trajectories $\mathbf{Z_{1}}(t)$ and $\mathbf{Z_{2}}(t)$ with initial separation vector $\delta\mathbf{Z}_{0}:=\mathbf{Z_{2}}(0)-\mathbf{Z_{1}}(0)$, a number $\lambda$ is called the Lyapunov exponent if it approximates the rate of divergence as $|\delta\mathbf{Z}(t)|\approx e^{\lambda t}|\delta\mathbf{Z_{0}}|$. Since this value may differ based on the orientation of initial separation vector, the maximal Lyapunov exponent is often used to characterize the behavior of dynamical systems and is defined as:
$\lambda=\lim_{t\rightarrow\infty}\lim_{|\delta\mathbf{Z_{0}}|\rightarrow 0}\frac{1}{t}\ln\frac{|\delta\mathbf{Z}(t)|}{|\delta\mathbf{Z}_{0}|}$
Positive values of $\lambda$ typically signify chaotic behavior in a system. For further background on the Lyapunov exponent, we refer the reader to Arnold & Wihstutz [2].
Chaos can be observed on dynamical systems on graphs. As a prototypical example, we consider the general, complex, nonlinear graph dynamics on graphs described in Sprott [37].
$\frac{d}{dt}u_{S}(v_{i},t)=-b\cdot u_{S}(v_{i},t)+\tanh\left(\sum_{v_{j}\in\mathcal{N}(v_{i})}a_{ij}u(v_{j},t)\right),\quad u_{S}(v,0)=\mathbf{x}(v)$ (5)
where $b$ is a damping coefficient. We refer to the dynamics governed by Eq. 5 as the Sprott equation and denote its solutions by $u_{S}(v,t)$. When $b>0$, solutions $u_{S}(v,t)$ remain bounded. In fully connected networks with $b=0.25$, positive Lyapunov exponents were observed when a sufficiently large fraction of interactions were neither symmetric ($a_{ij}=a_{ji}$) nor anti-symmetric ($a_{ij}=-a_{ji}$). Sparse connected networks also exhibited positive Lyapunov exponents at a value of $b=0.25$ [37].
3 Methods
3.1 Motivation: Connection between dynamics and topological properties
In this section, we formulate properties of the dynamics discussed above to establish their connection to graph topology and structure. These results serve as motivation for our method which utilizes these dynamics as a a novel message passing paradigms in a graph neural networks. For proofs, as well as more detailed statements when appropriate, please see Appendix A.
We begin by discussing a desirable property of solutions to all of the dynamics under consideration.
Proposition 3.1 (Connected components).
Let $u(v,t)$ denote either $u_{H}(v,t)$, $u_{W}(v,t)$, or $u_{S}(v,t)$ from Eqns. 3, 4, and 5 respectively. Suppose the non-zero entries of the initial condition $\mathbf{x}$ (and also $\mathbf{y}$ in the wave case) are contained in a given connected component. Then, the non-zero entries of $u(\cdot,t)$ are contained in that same component for all $t$.
Due to Proposition 3.1 we make the assumption that $G$ is connected in the following sections. However, we note that many of the results still apply to disconnected graphs with suitable modifications.
We now discuss several desirable attributes of the heat, wave, and Sprott dynamics which stem from the structure of the differential equation and their corresponding solutions. The dynamics we consider are canonical and capture a wide range of solution behavior, thereby enabling the results of this work to set the groundwork for future dynamics based message passing schemes.
3.1.1 Heat Dynamics
The continuous and global nature of DYMAG allows the model to instantaneously have a receptive field over the entire graph, thereby ameliorating the under-reaching problem in traditional message passing. The following proposition shows that $u(v,t)$ is strictly positive for all $t>0$.
Proposition 3.2.
Let $G$ be connected and let $L$ be given by the random-walk Laplacian $L_{rw}$. Assume that $\mathbf{x}(v)\geq 0$ for all $v$ (with at least one $\mathbf{x}(v)>0$), then $u_{H}(v,t)>0$ for all $v\in V$ and $t>0$.
Intuitively, this corresponds to information spreading instantaneously over the entire graph (although for small values of $t$ the energy $u(v,\cdot)$ will be mostly concentrated near the support of $\mathbf{x}$). This is in contrast to message-passing networks where the receptive field about each node is equal to the number of layers, which is typically kept small in order to avoid oversmoothing.
Proposition 3.3 (Heat energy).
Let $G$ be connected, and let $u_{H}(v,t)$ be as in Eqn. 3. Then,
$e^{-2t\lambda_{n}}\|\mathbf{x}\|_{2}^{2}\leq\|u_{H}(\cdot,t)\|_{2}^{2}\leq\langle\boldsymbol{\nu_{1}},\mathbf{x}\rangle^{2}+e^{-2t\lambda_{2}}\|\textbf{x}\|_{2}^{2}.$
This proposition suggests that graphs with a larger $\lambda_{2}$ will have a faster rate of energy decay. The second eigenvalue of a graph can be related to the isoperimetric ratio of a graph through Cheeger’s inequality, thereby revealing information on graph structure and how “bottlenecked" a particular graph is [36]. This property of heat energy decay can be extended to comparisons between graph structures, as illustrated by the following proposition:
Proposition 3.4 (Heat energy between graphs).
Let $G$ and $G^{\prime}$ be graphs on $n$ vertices with respective Laplacians $L_{G}$ and $L_{G^{\prime}}$ and let $\mathbf{x}$ and $\mathbf{x}^{\prime}$ be initial conditions for the heat equation (Eqn. 1) on $G$ and $G^{\prime}$. Assume that:
  1. 1.
    $L_{G^{\prime}}\succcurlyeq L_{G}$, i.e., $\mathbf{v}^{T}L_{G^{\prime}}\mathbf{v}\geq\mathbf{v}^{T}L_{G}\mathbf{v}\>\>\text{for all }\mathbf{v}\in\mathbb{R}^{n},$
  2. 2.
    We have $\langle\boldsymbol{\nu_{i}^{\prime}},\mathbf{x^{\prime}}\rangle^{2}\leq(1+\delta_{i}(t))\langle\boldsymbol{\nu_{i}},\mathbf{x}\rangle^{2}$ for all $1\leq i\leq n,$ where we also assume $\delta_{i}(t):=e^{2t(\lambda_{i}^{\prime}-\lambda_{i})}-1\geq 0.$
Then, with $u_{H}$ and $u_{H}^{\prime}$ are defined as in Eqn. 3, we have
$\|u_{H}^{\prime}(t)\|^{2}_{2}\leq\|u_{H}(t)\|^{2}.$ (6)
We note that Condition 1 of the proposition is satisfied if $G^{\prime}$ can be constructed from $G$ by the addition of edges or by increasing the weight of any edges [36]. Condition 2 may be satisfied by choosing $\mathbf{x}$ and $\mathbf{x}^{\prime}$ to be sum of all eigenvectors $\boldsymbol{\nu_{i}}$ and $\boldsymbol{\nu_{i}^{\prime}}$. Although the assumptions for Proposition 3.4 represents a rather specific set of conditions, we expect that when two graphs have edges generated according to a similar rule or distribution, the more densely connected graph will have more rapidly decaying heat energy. This is explored in Figure 1 with Erdos-Renyi random graphs.
Finally, we restate a known results that provide additional foundation linking the behavior of the heat equation solutions to graph topology. Lemma 1 of Crane et al. [12] shows that the heat equation encodes shortest path distances $d(u,v)$ between nodes $u,v$ on the graph:
Proposition 3.5 (Relation to distances, (Lemma 1 of Crane et al. [12])).
Let $u_{H}$ denote the solution to the heat equation with initial condition $\delta_{v}$. Then,
$d(v,w)=\lim_{t\to 0}\frac{\log\tilde{u}_{H}(w,t)}{\log t}.$ (7)
We additionally note that the ability of the heat equation to capture intrinsic graph distances links it to graph Ollivier-Ricci curvature $\kappa$ and provide a detailed statement in the appendix.
3.1.2 Wave Dynamics
The periodicity of the solution to the wave equation endows it with the ability to capture long range interactions. Central to this argument is the wave energy, analyzed in the following proposition:
Figure 1: $\|u_{H}(t)\|_{2}^{2}$ plotted for Erdos-Renyi graphs with 25 nodes and probability of connection between nodes given by $p$. Average results are reported over 30 Erdos-Renyi graphs for each $p$. For each graph a node is randomly selected to have a Dirac as the initial condition. It is observed that $\|u_{H}(t)\|_{2}^{2}$ decays more rapidly for Erdos-Renyi graphs with a larger $p$.
Proposition 3.6 (Wave energy).
Let $u_{W}(v,t)$ be the solution to the wave equation defined as in Eqn. 4 and assume that $\mathbf{y=0}$ in Eqn. 2. Then, for any $t\geq 0$, we have
$\langle\boldsymbol{\nu_{1}},\mathbf{x}\rangle^{2}\leq\|u_{W}(t)\|_{2}^{2}\leq\|\mathbf{x}\|_{2}^{2}=\|u_{W}(0)\|_{2}^{2}.$ (8)
Eqn. 8, shows that unlike the heat energy which is non-increasing, the wave energy will oscillate over time. Thus, the wave equation may be well suited for capturing global geometric features which require long-range interactions. The form that the solution to the wave equation takes on makes it particularly well suited to capturing aspects of graph topology. In particular, we find that it provides global information from local observation.
Proposition 3.7 (Recovery of eigenvalues).
The full Laplacian eigenspectrum $\lambda_{1},\ldots,\lambda_{n}$ can be approximated, up to an arbitrarily small degree of precision, from finitely many samples of the wave equation solution $u_{W}(v,t_{1}),u_{W}(v,t_{2}),\ldots,u_{W}(v,t_{m})$ at a single node $v$.
The graph spectra encodes a wide range of graph invariants and properties. The above proposition demonstrates that the solutions to the wave equation relate to graph spectral properties, and that this entire information is contained in the solutions of the wave equation at each node. As an example of how this may be related to graph topology, for instance in a cycle graph, the eigenvalues determine the length of the cycle. More generally, when the graph is not a cycle but contains a cyclic subgraphs as a prominent topological feature, this proposition provides some intuition for why the wave-equation is well suited to pick up that a node belongs to a cycle and recover dimension 1 homology.
3.1.3 Chaotic dynamics
Based on choice of $b=0.25$, it was observed in Sprott [37] that certain networks exhibited weakly chaotic (small positive Lyapunov exponents) dynamics. Weakly dynamic systems retain some memory but can explore much larger portions of the state space, thereby creating highly complex representations. In Lai & Winslow [24] a spatio-temporal chaotic dynamical system demonstrated highly sensitive dependence to initial condition in phase space and in the parameter space of this system. Thus chaotic network dynamics may be useful in situations where it is important to distinguish small differences in graph structure and data.
3.2 The DYMAG Architecture
Figure 2: The DYMAG architecture. Given an input signal on the graph, we first decompose it into $n$ Dirac signals. We then solve PDEs and collect the results using statistical moments, to obtain the node embedding. We can further collect the node embeddings to form a graph embedding. Depending on the task type, the appropriate embedding is fed into an MLP to get the model output.
The DYMAG architecture contains two main parts: The multiscale dynamics-based message passing scheme and the back-end neural network. The first part solves the relevant equations on the graph, and obtains a corresponding feature representation. The neural network takes that representation as input and is trained for the machine learning task.
The message passing consists of solving each equation with $n$ different initial conditions and using the solutions to generate the node representations $\mathbf{x}^{(i)},1\leq i\leq n$ (and also $\mathbf{y}^{(i)},1\leq i\leq n$ in the case of the wave equation). To construct these initial conditions, we assume that we are initially given a signal $\tilde{\mathbf{x}}\in\mathbb{R}^{n}$, (and also $\tilde{\mathbf{y}}\in\mathbb{R}^{n}$ if appropriate). For $1\leq i\leq n$, we define $\mathbf{x}^{(i)}=\tilde{\mathbf{x}}(v_{i})\boldsymbol{\delta}_{v_{i}},$ (and similarly $\mathbf{y}^{(i)}=\tilde{\mathbf{y}}(v_{i})\boldsymbol{\delta}_{v_{i}},$) and think of $\mathbf{x}^{(i)}$ and $\mathbf{y}^{(i)}$ as Dirac signals on the $i$-th node rescaled by the initial signal evaluated at $v_{i}$. If the data set does not have a signal $\tilde{\mathbf{y}}$, we set $\mathbf{y}^{(i)}=\mathbf{0}$ or $\mathbf{y}^{(i)}=\mathbf{x}^{(i)}$. Furthermore, if there is no initial signal in the dataset, we set $\mathbf{x}^{(i)}=\boldsymbol{\delta}_{v_{i}}$.
Let $u^{(i)}(v,t)$ denote the solution with the initial condition $u(\cdot,0)=\mathbf{x}^{(i)}$ (and $\partial_{t}u(\cdot,t)=\mathbf{y}^{(i)}$ if applicable). We collect these solutions to a more compact tensor using statistical moments. For fixed parameters $K$ and $M$, the hidden representation of node $v_{i}$ at time $t$ is given by
$h_{i}(t)=\Bigg{\{}\Big{\{}\sum_{j\in\mathcal{N}_{k}(i)}|u^{(i)}(v_{j},t)|^{m}\Big{\}}_{m=1}^{M}\Bigg{\}}_{k=1}^{K},$ (9)
where $\mathcal{N}_{k}(i)$ denotes the set of vertices within $k$ hops of $v_{i}$. Evaluating $h_{i}(t)$ at $t=1,2,\ldots,T$, we obtain a $M\times K\times T$ dimensional hidden representation for vertex $v_{i}$. (Our method can also be extended to the case where there are $r>1$ initial signals, in which case we simply repeat this procedure $r$ times and then concatenate.) For graph-level tasks, we compute statistical moments up to some fixed order $S$ of the representations along the axis of the nodes as a means of pooling:
$w(t)=\Big{\{}\sum_{i=1}^{n}|h_{i}(t)|^{s}\Big{\}}_{s=1}^{S}.$ (10)
Depending on if the task is defined at the node or graph level, we pass a flattened version of the node level representation (Eqn. 9) or the graph level representation (Eqn. 10) to a multi-layer perceptron (MLP). This is the trainable part of DYMAG. The construction of DYMAG is flexible and amenable to incorporation with other models. We leave such exploration to future work. Our code is publicly available [^1] https://anonymous.4open.science/r/DYMAG-196E/, and additional implementation details are described in Section B of the appendix.
Notably, the time dimension of the representation generated by DYMAG is unique when compared to other graph neural network methods. This choice is motivated by theoretical results discussed in Section 3.1 and allows DYMAG to infer certain topological properties of the graph.
4 Empirical Results
We conduct a wide range of experiments to assess the ability of various methods to recover aspects of graph structure and topology. We compare against a traditional message passing GNN, Graph Attention Network (GAT) [41], Graph Isomorphism Network (GIN) [43], Graph Neural Diffusion (GRAND) [8], and GRAND++ [39]. When recovering the extended persistence image we compare against Persistence Diagram Graph Neural Network (PDGNN) [44], a purpose-built graph neural network with customized pooling and message passing operations to approximate the Union-Find algorithm for predicting persistent homology. In our tables, best and second best are highlighted.
Table 1: Mean squared error (MSE) for prediction of extended persistence image with vertex degree as node features (lower is better). The first group of three ER graphs are created with $n=100$ nodes and the second is created using $n=200$ nodes. A full version of the table with standard deviation is provided in Section C.2 of the appendix.
Graph Extended Persistence Image
DYMAG PDGNN GNN GAT GIN GRAND GRAND++
(Heat) (Wave) (Sprott)
$\text{ER}(p=0.04)$ 1.48e-2 6.37e-3 6.63e-1 3.07e-3 7.83e-1 3.82 4.83e-1 7.72e-1 2.39
$\text{ER}(p=0.06)$ 8.65e-3 2.79e-3 6.24e-1 1.37e-3 7.35e-1 1.57 4.09e-1 5.72e-1 1.15
$\text{ER}(p=0.08)$ 8.82e-3 2.54e-3 5.71e-1 1.90e-3 9.46e-1 6.65e-1 3.92e-2 1.07e-2 8.59e-1
$\text{ER}(p=0.04)$ 8.91e-3 5.18e-3 8.04e-2 3.24e-3 6.73e-1 1.93e-1 3.72e-2 7.31e-2 2.85e-1
$\text{ER}(p=0.06)$ 7.41e-3 4.76e-3 1.39e-1 4.30e-3 7.29e-1 5.48e-1 3.69e-2 8.39e-2 3.28e-1
$\text{ER}(p=0.08)$ 4.57e-3 1.62e-3 1.35e-1 1.12e-3 1.28 1.87e-1 3.96e-2 4.12e-2 3.48e-1
Cora 1.45e-4 7.41e-5 3.51e-3 5.80e-5 2.72e-3 5.04e-2 5.53e-3 8.13e-4 8.16e-3
Citeseer 3.94e-4 1.58e-4 7.64e-3 1.29e-4 6.83e-3 2.93e-2 3.94e-3 6.87e-4 4.21e-3
PubMed 7.93e-3 3.29e-3 8.96e-2 2.74e-3 4.29e-1 4.79e-1 1.27e-1 3.37e-2 3.07e-1
Table 2: Mean squared error (MSE) for the prediction of generating parameters of random graphs (lower is better). The number of nodes for each type of random graph is specified in each data column (i.e. $n\in\{100,250,500,1000,2500\}$).
Method Erdős-Rényi Stochastic Block Model
$100$ $250$ $500$ $1000$ $2500$ $100$ $250$ $500$ $1000$ $2500$
DYMAG (Heat) 7.46e-3 7.13e-3 3.60e-3 4.19e-3 3.04e-3 6.41e-1 8.10e-1 1.79 4.52 11.63
DYMAG (Wave) 8.29e-3 6.58e-3 3.17e-3 3.25e-3 1.04e-3 8.25e-1 9.40e-1 1.26 2.28 2.35
DYMAG (Sprott) 4.33e-2 4.92e-2 7.08e-3 3.68e-3 5.49e-3 5.17 3.37 4.25 4.08 6.27
GNN 1.37e-2 1.14e-2 9.26e-3 9.49e-3 8.02e-3 2.93 3.07 3.68 7.14 10.26
GAT 3.05e-2 5.60e-2 1.35e-2 3.74e-2 2.69e-2 11.79 9.42 10.83 13.62 18.60
GIN 1.08e-2 9.37e-3 7.74e-3 6.98e-3 4.81e-3 1.74 2.59 2.92 4.37 9.15
GRAND 6.36e-2 4.22e-2 9.27e-3 6.58e-3 5.30e-3 14.52 16.78 13.50 11.28 8.58
GRAND++ 8.52e-2 6.91e-2 2.84e-2 1.29e-2 8.72e-3 23.71 26.84 19.64 16.97 15.42
Table 3: R2 score (higher is better) of DYMAG trained to predict geometric and graph topological properties on the PROTEINS [15] and DrugBank [42] datasets. Mean and standard deviation are calculated using $10$-fold cross-validation.
Model PROTEINS ($1,113$ proteins) DrugBank ($12,416$ molecules)
Dihedral Angles Bond Length TPSA # Aromatic Rings
DYMAG (Heat) ${\color[rgb]{0,0,1}\mathbf{0.87\pm 0.03}}$ ${\color[rgb]{.75,0,.25}\mathbf{0.96\pm 0.03}}$ ${\color[rgb]{0,0,1}\mathbf{0.92\pm 0.01}}$ ${\color[rgb]{0,0,1}\mathbf{0.96\pm 0.02}}$
DYMAG (Wave) $0.74\pm 0.06$ ${\color[rgb]{0,0,1}\mathbf{0.98\pm 0.01}}$ ${\color[rgb]{.75,0,.25}\mathbf{0.86\pm 0.02}}$ ${\color[rgb]{.75,0,.25}\mathbf{0.85\pm 0.02}}$
DYMAG (Sprott) ${\color[rgb]{.75,0,.25}\mathbf{0.78\pm 0.02}}$ $0.76\pm 0.03$ $0.73\pm 0.04$ $0.79\pm 0.02$
GNN ${\color[rgb]{.75,0,.25}\mathbf{0.78\pm 0.01}}$ $0.89\pm 0.03$ $0.71\pm 0.01$ $0.81\pm 0.01$
GAT $0.72\pm 0.02$ $0.91\pm 0.01$ $0.78\pm 0.02$ $0.83\pm 0.02$
GIN $0.69\pm 0.03$ $0.93\pm 0.01$ $0.69\pm 0.01$ $0.77\pm 0.01$
GRAND $0.76\pm 0.03$ $0.87\pm 0.02$ $0.53\pm 0.04$ $0.64\pm 0.03$
GRAND++ $0.62\pm 0.03$ $0.73\pm 0.03$ $0.56\pm 0.02$ $0.61\pm 0.02$
4.1 Prediction of geometrical and topological properties of graphs
First, we assessed the expressivity of multiscale dynamics as replacements for message passing on graphs by training our network to recover geometric and topological properties. We trained the network to predict the persistence image representation [1] of the extended persistence diagram. The extended persistence diagram is obtained from ascending and descending filtrations of the $5$-hop neighborhood graphs at each node, using node degree as the filter function [11, 44]. The extended persistence image consists of connected components (dimension 00 homology) and topological loops (dimension $1$ homology).
We trained and tested the model on Erdős-Rényi graphs, $G(n,p)$, with $n=\{100,200\}$ vertices and $p=\{0.04,0.06,0.08\}$ edge probabilities as well as various citation graphs. We used initial conditions with a Dirac (Kronecker) $\boldsymbol{\delta}$ at each node and computing the moments of the solution across the nodes at various timepoints.
Our results demonstrate that using dynamics in lieu of traditional message-passing results in significantly higher prediction accuracy on persistent homology of graphs when compared to a variety of message-passing GNNs and GRAND and Table 1). The performance of DYMAG is on par with PDGNN despite not being explicitly designed for this task. The wave dynamics were best for DYMAG, and our network showed good generalization capabilities across different graph structures, across both synthetic and real-world datasets, including Cora [25], Citeseer [19], and PubMed [34]. Additional experiments on recovering the extended persistence image with a different filtration function are presented in Section C.2 of the appendix.
Next, we considered the task of recovering the structure of random graphs generated using the Erdős-Rényi (ER) model and the stochastic block model (SBM). Specifically, we aim to recover the generative parameters for these models, namely the probability of connecting two nodes by an edge in the ER model and the number of blocks/communities in the SBM. DYMAG outperformed message passing networks and GRAND in this task (Table 2), demonstrating the expressivity of PDE solutions and their ability to identify the generative process underlying random graph construction.
Additional experiments in Section C.1 of the appendix demonstrate the superior ability of DYMAG to recover the Ollivier-Ricci curvature of the graph nodes.
4.2 Molecular classification and property prediction
We next conducted experiments with molecular graphs representing proteins and drug-like molecules (Tables 3 and 4). The geometry and topology of proteins are defined by the arrangement of their amino acid residues in the three-dimensional space. In particular, the overall shape of the protein is defined by dihedral angles, which measure the rotation of the amino acid residues around the bonds, and bond lengths between successive amino acids in the protein sequence. As shown in Table 3, DYMAG outperforms GNN, GAT, GIN, GRAND, and GRAND++ in predicting the dihedral angles and bond lengths of proteins in the PROTEINS [15] dataset.
We also compared DYMAG with other graph neural networks in predicting two important properties that influence their biological activity and pharmacokinetics of drug-like molecules. The total polar surface area (TPSA) is the surface area occupied by polar atoms (oxygen, nitrogen, and attached hydrogens) in a molecule. TPSA is related to the ability of a molecule to cross biological membranes, like the blood-brain barrier, and is therefore, an important factor in drug design. As shown in Table 3, DYMAG trained using the wave equation on molecular graphs, outperforms standard message-passing GNN, GAT, and GIN in predicting TPSA and number of aromatic rings in drug-like molecules.
In general, the heat and wave versions of DYMAG had excellent performance on all tasks while the Sprott version produced mixed results. This highlights a limitation of our method: many tasks are not well suited for chaotic dynamics where high sensitivity to graph signal and structure is undesirable. We leave exploration of such systems to future work.
5 Related work
There is a long history of studying dynamics on graphs. For example, Reijneveld et al. [31], Boccaletti et al. [4], Simoes [35], Holme & Saramäki [21] analyze complex interactions such as brain processes, social networks, and spatial epidemics on graphs. The study of graph dynamics has recently crossed into the field of deep learning. with Belbute-Peres et al. [3], Sanchez-Gonzalez et al. [32], Pfaff et al. [30] using neural networks to simulate complex phenomena on irregularly structured domains.
Here, differing from the above mentioned works, we aim to understand GNNs and their ability to capture topological properties. Using dynamics to understand GNNs has been the subject of several recent works. Chamberlain et al. [7] and Chamberlain et al. [8] take the perspective that message-passing neural networks can be interpreted as the discretizations of diffusion-type (parabolic) partial differential equations on graph domains where each layer corresponds to a discrete time step. They then use this insight to design GRAND, a novel GNN, based on encoding the input node features, running a diffusion process for $T$ seconds and finally applying a decoder network. Thorpe et al. [39] builds on this work by extending it to diffusion equations with “sources” placed at the labeled nodes, leading to a new network GRAND++. They then provide an analysis of both GRAND and GRAND++ and show that they are related to different graph random walks. GRAND is related to a standard graph random walk, whereas GRAND++ is related to a dual random walk started at the the labeled data which can avoid the oversmoothing problem. Moreover, we note Donnat et al. [16], which used the solution to of the graph heat equation to extract structural information around each node (although not in a neural network context) and Eliasof et al. [17] which used insights from hyperbolic and parabolic PDEs on manifolds to design a GNN that does not suffer from oversmoothing.
Our network method differs from these previous works in several important ways. Whereas Chamberlain et al. [8] and Chamberlain et al. [7] primarily focused on PDEs as a framework for understanding the behavior of message passing operations, here we propose to use the dynamics associated to the heat equation as a new form of continuous-time message passing, replacing traditional discrete-time message passing operations. We consider both the heat equation (the protypical parabolic PDE) and the wave equation (the prototypical hyperbolic PDE) as well as chaotic dynamics, whereas previous work [7, 8, 39] has primarily focused on parabolic equations. Notably, similar to GRAND++, the wave-equation and the Sprott versions of DYMAG do not suffer from oversmoothing. However, the long-term behavior of these equations differ from the diffusion-with-a-source equation used in GRAND++ in that they only depends on the geometry of the network and not on locations of the labeled data. (Additionally, since we does not require labeled data as source locations, our method can be easily adapted to unsupervised problems.) Finally, we note that we use a different method for extracting initial features and of performing statistical moment aggregation.
6 Conclusion, Limitation and Future Research
We have introduced DYMAG, a neural network which utilizes dynamics of heat and wave propagation to infer geometric and topological properties of a network. Our method is based upon (i) solving a PDE on the graph, (ii) extracting statistical moments from the solution at multiple time points, and (iii) training a multi-layer perceptron on top of these statistical moments. We provide a theoretical analysis showing that these dynamics are related to the connectivity of the graph, several notions of curvature, as well as other properties such as cycle lengths. In our numerical experiments, we demonstrate that our method is effective for predicting these topological properties as well as the generating parameters of random graphs drawn from different families.
DYMAG is not applicable to all tasks. In tasks such as node classification on homophilic networks, multiscale features, graph topology, highly discriminative features, and complex dynamics are not essential. In these cases simpler models should be preferred. Our future work seeks to consider additional dynamics and understanding of their ability to capture structural information.
Impact Statement
Graph representation learning that can capture structure is applicable to many domains. We expect our work to accelerate scientific discovery, particularly in fields such as biochemistry and neuroscience, where connectivity structure (between molecules, or neurons) plays an essential role. It is more generally applicable to any problem where graph structure in important to capture in a representation. Thus we see it as capable of having a positive effect in many areas of societal concern.
Appendix A Proofs of Theoretical Results
A.1 The proof of Proposition 3.1
The following is a more detailed statement of Proposition 3.1.
Proposition A.1 (Connected components).
Suppose $G$ has $m$ connected components and partition the vertex set $V=V_{1}\sqcup V_{2}\sqcup\ldots\sqcup V_{m}$ where each $V_{i}$ is a connected component. Let $S$ be a set that is obtained by taking the union of several $V_{i}$’s, i.e., $S=V_{i_{1}}\sqcup\ldots\sqcup V_{i_{k}}$. Let $u(v,t)$ denote either $u_{H}(v,t)$, $u_{W}(v,t)$ or $u_{S}(v,t)$, the solutions to the heat, wave, or chaotic dynamics equations. Assume that the if the initial condition $\mathbf{x}=u(\cdot,0)$ (and also $\mathbf{y}=\partial_{t}u(\cdot,0))$ in the wave case) has support contained in $S$. Then for any $v\notin S$ and for all $t>0$, we have
$u(v,t)=0.$ (11)
Proof.
For notational simplicity, assume $m=2,$ $S=V_{1}$. (The proof of the general case is identical.) Also observe it suffies to prove the result on for $0\leq t\leq T$ where $T$ is arbitary (since we may then let $T\rightarrow\infty$).
Let $u$ be a solution to the differential equation and consider the function $\tilde{u}$ defined by
$\tilde{u}(v,t)=\begin{cases}u(v,t)\quad&\text{if }v\in V_{1}\\ 0\quad&\text{if }v\in V_{2}\end{cases}.$
We observe that $\tilde{u}$ is also a solution since the right hand side of each differential equation is localized in the sense that no energy passes between components and the initial condition has support contained in $V_{1}$. However, since the right hand side of all three differential equations is Lipschitz on $[0,T]$, Theorem 5, Section 6 of Coddington [10] implies that there is at most one solution to the differential equation and thus $\tilde{u}=u$. Therefore, we have $u(v,t)=0$ for all $v\notin V_{1}$.
A.2 The Proof of Proposition 3.2
In order to prove Propostion 3.2, we need the following theorem result which relates the heat equation to a continuous time random walker [18] $\{X^{\text{continuous}}_{t}\}_{t\geq 0}$ defined by $X^{\text{continuous}}_{t}=X^{\text{discrete}}_{N(t)},$ where $\{X^{\text{discrete}}_{i}\}_{i=0}^{\infty}$ is a standard discrete-time random walker (i.e., a walker who moves one at each discrete time step) and $\{N(t)\}_{t\geq 0}$ is an ordinary Poisson process. For a proof, please see Section A.7.
Proposition A.2.
Let $u_{H}(v,t)$ be the solution to the heat equation with $L$ chosen to be the random-walk Laplacian, $L=L_{RW}=I-P$ where $P=D^{-1}A$ and initial condition $\mathbf{x}$. Assume that $\mathbf{x}$ is entrywise nonnegative with entries summing to one. Then $u(\cdot,t)$ is the probability distribution of a continuous-time random walker, with initial distribuion $\mathbf{x}$, at time $t$.
We may now prove Propostion 3.2.
Proof.
By linearity, it suffices to consider the case where the entries of $\mathbf{x}$ sum to one, in which case $\mathbf{x}$ may be interpretted as a probability distribution. Let $v_{1}\in V$ be arbitrary. Choose $v_{0}\in V$ such that $\mathbf{x}(v_{0})>0$. As shown in Theorem A.2, $u(\cdot,t)$ is the probability distribution of a continuous-time random walk with initial distribution $\mathbf{x}$ at time $t$. Therefore, if $d$ is the length of the shortest path from $v_{0}$ to $v_{1}$, then
$u(v_{1},t)$ $=\mathbb{P}(X^{\text{continuous}}_{t}=v_{1}|X^{\text{continuous}}_{0}\sim\mathbf{x})$
$\geq\mathbb{P}(N(t)=d)\mathbb{P}(X_{d}=v_{1}|X_{0}=v_{0})>0.\qed$
A.3 The Proof of Proposition 3.3
The following is a more detailed statement of Proposition 3.3.
Proposition A.3 (Heat energy).
Let $G$ be connected, and let $u_{H}(v,t)$ be as in Eqn. 3 and let $u_{H}(t)=u_{H}(\cdot,t)$. Then,
$e^{-2t\lambda_{n}}\|\mathbf{x}\|_{2}^{2}\leq\|u_{H}(t)\|_{2}^{2}\leq\langle\boldsymbol{\nu_{1}},\mathbf{x}\rangle^{2}+e^{-2t\lambda_{2}}\|\textbf{x}\|_{2}^{2}$ (12)
for all $t>0$. The upper bound on heat energy obeys:
$\|u_{H}(t)\|_{2}^{2}\leq\|\mathbf{x}\|_{2}^{2}=\|u_{H}(0)\|_{2}^{2}.$ (13)
Additionally, as $t$ converges to infinity we have
$\lim_{t\rightarrow\infty}u_{H}(t)=\langle\boldsymbol{\nu_{1}},\mathbf{x}\rangle\boldsymbol{\nu_{1}}.$ (14)
Proof.
From Eqn. 3, and the fact that $\{\boldsymbol{\nu_{i}}\}_{i=1}^{n}$ is an ONB, we see
$\|u_{H}(t)\|_{2}^{2}$ $=\left\langle\sum_{i=1}^{n}e^{-t\lambda_{i}}\langle\boldsymbol{\nu_{i}},\mathbf{x}\rangle\boldsymbol{\nu_{i}},\sum_{i=1}^{n}e^{-t\lambda_{i}}\langle\boldsymbol{\nu_{i}},\mathbf{x}\rangle\boldsymbol{\nu_{i}}\right\rangle$
$=\sum_{i=1}^{n}e^{-2t\lambda_{i}}\langle\boldsymbol{\nu_{i}},\mathbf{x}\rangle^{2}.$ (15)
Thus, since $\lambda_{1}=0$, upper bound in (12) is obtained by:
$\|u_{H}(t)\|_{2}^{2}$ $=\sum_{i=1}^{n}e^{-2t\lambda_{i}}\langle\boldsymbol{\nu_{i}},\mathbf{x}\rangle^{2}$
$=\langle\boldsymbol{\nu_{1}},\mathbf{x}\rangle^{2}+\sum_{i=2}^{n}e^{-2t\lambda_{i}}\langle\boldsymbol{\nu_{i}},\mathbf{x}\rangle^{2}$
$\leq\langle\boldsymbol{\nu_{1}},\mathbf{x}\rangle^{2}+e^{-2t\lambda_{2}}\sum_{i=2}^{n}\langle\boldsymbol{\nu_{i}},\mathbf{x}\rangle^{2}$
$\leq\langle\boldsymbol{\nu_{1}},\mathbf{x}\rangle^{2}+e^{-2t\lambda_{2}}\|\mathbf{x}\|_{2}^{2}.$
The first inequality in Eqn. 13 follows by noting:
$\sum_{i=1}^{n}e^{-2t\lambda_{i}}\langle\boldsymbol{\nu_{i}},\mathbf{x}\rangle^{2}$ $\leq\sum_{i=1}^{n}\langle\boldsymbol{\nu_{i}},\mathbf{x}\rangle^{2}=\|\mathbf{x}\|^{2}_{2}$
and the second follows by observing that equality holds when $t=0$. The lower bound in (12) may be obtained by noting
$\sum_{i=1}^{n}e^{-2t\lambda_{i}}\langle\boldsymbol{\nu_{i}},\mathbf{x}\rangle^{2}$ $\geq e^{-2t\lambda_{n}}\sum_{i=1}^{n}\langle\boldsymbol{\nu_{i}},\mathbf{x}\rangle^{2}=e^{-2t\lambda_{n}}\|\mathbf{x}\|^{2}_{2}.$
Eqn. 14 immediately follows Eqn. 15 and the fact that $0=\lambda_{1}<\lambda_{2}\leq\ldots\leq\lambda_{n}$. ∎
A.4 The Proof of Proposition 3.4
Proof.
By (15), we have
$\|u_{H}(t)\|_{2}^{2}-\|u_{H}^{\prime}(t)\|^{2}_{2}=$ $\sum_{i=1}^{n}e^{-2t\lambda_{i}}\langle\boldsymbol{\nu_{i}},\mathbf{x}\rangle^{2}-e^{-2t\lambda_{i}^{\prime}}\langle\boldsymbol{\nu_{i}^{\prime}},\mathbf{x^{\prime}}\rangle^{2}$
$\geq\sum_{i=1}^{n}\left[e^{-2t\lambda_{i}}-e^{-2t\lambda_{i}^{\prime}}(1+\delta_{i})\right]\langle\boldsymbol{\nu_{i}},\mathbf{x}\rangle^{2}$
$=0.\qed$
A.5 Proof of Proposition 3.6
Proof.
The proof for Eqn. 8 is similar to the proof of Eqn. 12.
$\|u_{W}(t)\|_{2}^{2}$ $=\sum_{i=1}^{n}\cos^{2}(\sqrt{\lambda_{i}}t)\langle\boldsymbol{\nu}_{i},\mathbf{x}\rangle^{2}.$ (16)
Therefore,
$\|u_{W}(t)\|_{2}^{2}$ $=\sum_{i=1}^{n}\cos^{2}(\sqrt{\lambda_{i}}t)\langle\boldsymbol{\nu}_{i},\mathbf{x}\rangle^{2}\leq\sum_{i=1}^{n}\langle\boldsymbol{\nu}_{i},\mathbf{x}\rangle^{2}=\|\mathbf{x}\|_{2}^{2}.$
The lower bound follows by noting that since $\lambda_{1}=0$ we have:
$\|u_{W}(t)\|_{2}^{2}$ $=\sum_{i=1}^{n}\cos^{2}(\sqrt{\lambda_{i}}t)\langle\boldsymbol{\nu}_{i},\mathbf{x}\rangle^{2}$
$\geq\cos^{2}(\sqrt{\lambda_{1}}t)\langle\boldsymbol{\nu}_{1},\mathbf{x}\rangle^{2}$
$=\langle\boldsymbol{\nu}_{1},\mathbf{x}\rangle^{2}.\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qed$
A.6 The proof of Proposition 3.7
Proof.
Fix $v$. Since $\mathbf{y}=0,$ we may rewrite (4) as
$f(t)\coloneqq u_{W}(v,t)=\sum_{i=1}^{n}\cos(\sqrt{\lambda_{i}}t)c_{i}(v)$
where $c_{i}(v)=\langle\boldsymbol{\nu}_{i},\mathbf{x}\rangle\boldsymbol{\nu}_{i}(v)$ is a constant with respect to time and depends only on the node position.
Now, let $\epsilon>0$ be a degree of percision and choose $K$ such that $\frac{1}{K}<\epsilon$. Approximate
$f(t)\approx\tilde{f}(t)\coloneqq\sum_{i=1}^{n}\cos(a_{i}t)c_{i}(v)$
where $a_{i}$ is the multiple of $1/K$ such that $|a_{i}-\sqrt{\lambda_{i}}|<\epsilon$. The function $\tilde{f}$ has a finite Fourier expansion and therefore is uniquely characters by finitely many samples which allows us to recover the $a_{i}$ and thus approximately recover the $\lambda_{i}$.
A.7 The proof of Proposition A.2
Proof.
We first note that $L=L_{rw}$ can be written in terms of the symmetric normalized Laplacian $L_{s}=I_{n}-D^{-1/2}AD^{-1/2}$ as $L_{rw}=D^{-1/2}L_{s}D^{1/2}$. Therefore, $L_{rw}$ is diagonalizable and may be written as $L_{rw}=S\Lambda S^{-1}$ where $L_{s}=\Psi\Lambda\Psi^{-1}$ is a diagonalization of $L_{s}$ and $S=D^{-1/2}\Psi$. This allows us to write
$P=I-L_{rw}=S(I-\Lambda)S^{-1},$
which implies that for $k\geq 0$, we have
$P^{k}=S(I-\Lambda)^{k}S^{-1}.$
We next note that Eqn. 3 may be written as $u_{H}(\cdot,t)=U^{t}_{H}\mathbf{x}$
We define the heat kernel on the graph as $u_{H}(\cdot,t)=U^{t}_{H}\mathbf{x}$, where $U^{t}_{H}$ is the heat kernel defined by
$U^{t}_{H}:=e^{-tL}=Se^{-t\Lambda}S^{-1}=\sum_{i=1}^{n}e^{-t\lambda_{i}}\boldsymbol{\nu_{i}}(v)\boldsymbol{\nu_{i}}^{T}.$ (17)
Therefore, it suffices to show that $U^{t}_{H}$ is the $t$-second transition matrix of the continuous-time random walker $X^{\text{continuous}}_{t}=X^{\text{discrete}}_{N(t)}$.
By the definition of a Poisson process, $N(t)$ is a Poisson random variable with parameter $t$. Thus, for $k\geq 0,$, $t\geq 0$, we have
$\mathbb{P}(N_{t}=k)=A_{t}(k)=t^{k}e^{-t}/k!$
. We next observe that for all $\mu\in\mathbb{R}$ we have
$e^{-t(1-\mu)}=e^{-t}\sum_{k\geq 0}\frac{(t\mu)^{k}}{k!}=\sum_{k\geq 0}A_{t}(k)\mu^{k}.$ (18)
Substituting $\lambda=1-\mu$ in Eqn. 18 links the eigenvalues of $U^{t}_{H}$ and $P$ by
$U^{t}_{H}$ $=Se^{-t\Lambda}S^{-1}=S\sum_{k\geq 0}A_{t}(k)(I_{n}-\Lambda)^{k}S^{-1}$
$=\sum_{k\geq 0}A_{t}(k)P^{k}.$
This implies that $U^{t}_{H}$ is the $t$-second transition matrix of the continuous-time random walker and thus completes the proof.
A.8 Relation of the heat equation to Ollivier-Ricci Curvature
The Ollivier-Ricci curvature is a discrete notion of curvature, meant to paralell the tradition notion of Ricci curvature in Riemmanian geometry. It is defined by
$\kappa(v,w)=1-W_{1}(\mu_{v},\mu_{w})/d(v,w),$
where $\mu_{v}$ is the probability measure of the random walk at $u$, $W_{1}$ the 1-Wasserstein distance, and $d(v,w)$ is the distance (shortest path length) from $v$ to $w$.
The following result from Münch & Wojciechowski [27] related $\kappa$ to the solution of the heat equation.
Proposition A.4 (Relation to Ollivier-Ricci curvature, (Theorem 5.8 of [27])).
Let $L=D-A$ be the unnormalized Laplacian, then the Ollivier Ricci curvature $\kappa$ can be expressed as
$\kappa(v,w)=\lim_{t\to 0^{+}}\frac{1}{t}\bigg{(}1-\frac{W_{1}(u^{v}_{H}(\cdot,t),u^{w}_{H}(\cdot,t))}{d(v,w)}\bigg{)}$ (19)
where, for $v\in V$, $\delta_{v}$ is the Kronecker (Dirac) delta on the node $v$, $u_{H}^{v}(\cdot,\cdot)$ is the solution to the heat equation with initial condition $\delta_{v}$, and $W_{1}$ is the 1-Wasserstein distance.
Appendix B Implementation Details
B.1 Experimental Computation Resources
Throughout the experiments, we use a server with CPU type 6240 with 36 nodes (730 gigabytes) and GPU type a100 with 4 GPUs (40 gigbytes vRAM/GPUs).
B.2 DYMAG Parameters
Here we describe the parameter values of DYMAG described in Section 3.2 used to generate the experimental results.
Fixed values of $M=4$, $K=5$, and $S=3$ were used for all dynamics and models. $T$ was selected using grid search for each dataset. We used a 5-layer MLP with LeakyReLU activation to transform the node features obtained from heat and wave equation solutions to the 25-dimensional persistence image. The dimensionality of the hidden layers, in order from the input layer to the output layer, are $64$, $48$, and $32$.
B.3 DYMAG Dynamics Solver Details
Our implementation makes use of a Chebyshev polynomial approximation to compute the solutions to the heat and wave equations defined in Eqns. 3 and 4 respectively. In doing so we avoid eigendecomposing the Laplacian, which can have a computational complexity of $O(n^{3})$ and require $O(n^{2})$ memory [20]. The polynomial approximation has linear complexity for sparse graphs [14]. For example, if $G$ is a $k$-nearest neighbour graph and the order of the polynomial is $m$, then the time complexity for solving the heat/wave equation is $O(kmn)$. For the Sprott dynamics, a closed form solution is not available. Instead, we solve the equations using a simple forward Euler numerical method.
Our solutions to the graph heat and wave equations require Chebyshev approximations of the functions $e^{-tx},\cos(tx),$ and $\sin(tx)/(tx)$. Each of these functions can be extended to an entire function on the complex plane. Therefore, it follows from known results in approximation theory (see, e.g., Theorem 8.2 of Trefethen [40] with $\rho=2$), together with Parseval’s theorem, that
$\|u_{H}^{(K)}(\cdot,t)-u_{H}(\cdot,t)\|_{2}\leq\mathcal{O}(2^{-K})\|\mathbf{x}\|_{2}.$
where $u_{H}(v,t)$ is the solution to the heat equation (3) with initial value $\mathbf{x}$, $u^{(K)}_{H}(v,t)$ is its approximation by a $K$-th order Chebyshev approximation, and that an analogous inequality holds for the wave equation, but on the right-hand side, $\|\mathbf{x}\|_{2}$ is replaced by $\max\{\|\mathbf{x}\|_{2},\|\mathbf{y}\|_{2}\}$. We note that Perlmutter et al. [29] conducted an ablation study where it was found that the use of Chebyshev polynomials for approximation in a GNN-like architecture did not significantly change the accuracy.
Appendix C Additional Experiments
In this section, we provide additonal experimental validation of DYMAG concerning its ability to predict properties of biomolecules as well as topological properties such as curvature.
Table 4: Classification accuracy and runtime of DYMAG on ENZYMES [33, 5], PROTEINS [15, 5], and MUTAG [13, 23] datasets taken from the TUDatasets benchmarks collection [26]
Dataset DYMAG Model Training Accuracy Validation Accuracy Training Time (ms)
ENZYMES Heat ${\color[rgb]{0,0,1}\mathbf{0.85\pm 0.10}}$ ${\color[rgb]{0,0,1}\mathbf{0.82\pm 0.17}}$ 551 $\pm$ 4
Wave $0.80\pm 0.14$ $0.79\pm 0.21$ $400\pm 3$
PROTEINS Heat ${\color[rgb]{0,0,1}\mathbf{0.79\pm 0.02}}$ $0.54\pm 0.04$ $1045\pm 13$
Wave ${\color[rgb]{0,0,1}\mathbf{0.78\pm 0.02}}$ ${\color[rgb]{0,0,1}\mathbf{0.71\pm 0.07}}$ ${\color[rgb]{0,0,1}\mathbf{857\pm 11}}$
MUTAG Heat $0.86\pm 0.03$ $0.79\pm 0.09$ 219 $\pm$ 1
Wave ${\color[rgb]{0,0,1}\mathbf{0.87\pm 0.04}}$ ${\color[rgb]{0,0,1}\mathbf{0.83\pm 0.07}}$ $591\pm 3$
We evaluated the performance of DYMAG on three publicly available datasets for biomolecular graph classification from the TUDatasets benchmark [26]. The ENZYMES dataset [33] consists of protein secondary structures with ground truth annotations of catalytic activity. In the PROTEINS dataset [15], the task is to classify whether a protein functions as an enzyme. The MUTAG dataset [13], contains small nitroaromatic compounds, and the task is to classifiy their mutagenicity on the S.typhimurium bacterium. The DYMAG achieved high validation accuracy on these models, and the wave equation performed better on PROTEINS and MUTAG.
C.1 Ollivier-Ricci Curvature
We additionally perform experiments regressing the Ollivier-Ricci curvature [28] of a graph. Ricci curvature is used in Riemannian geometry to measure the deviation in the manifold from being locally Euclidean along the tangent directions. In the discrete graph setting, Ollivier-Ricci curvature captures the local geometry of the graph.
We used a 5-layer MLP to estimate Ollivier-Ricci curvature based on the node representations (Equation 9).
Table 5: Mean squared error (MSE) for prediction of curvature with vertex degree as node features (lower is better). The first group of three Erdös-Renyi (ER) graphs are created with $n=100$ nodes and the second is created using $n=200$ nodes.
Graph Ollivier-Ricci Curvature
DYMAG GNN GAT GIN GRAND GRAND++
(Heat) (Wave) (Sprott)
$\text{ER}(p=0.04)$ 1.86e-1 1.93e-1 7.44 32.04 23.70 5.93 11.04 14.85
$\text{ER}(p=0.06)$ 1.80e-1 1.76e-1 7.39 31.93 21.26 2.06 9.26 16.13
$\text{ER}(p=0.08)$ 1.78e-1 1.79e-1 6.81 31.87 29.91 8.62e-1 2.27 13.48
$\text{ER}(p=0.04)$ 3.63e-1 3.58e-1 1.52 18.09 6.74 7.86e-1 5.90e-1 2.06
$\text{ER}(p=0.06)$ 3.19e-1 2.63e-1 5.28e-1 17.43 4.39 3.39e-1 7.38e-1 1.82
$\text{ER}(p=0.08)$ 2.14e-1 2.57e-1 5.73e-1 15.50 6.18 4.27e-1 4.33e-1 1.76
Cora 1.73e-2 1.85e-2 1.34e-1 2.40e-1 7.36e-1 1.56e-1 6.81e-2 1.74e-1
Citeseer 2.09e-2 3.41e-2 1.46e-1 2.20e-1 9.04e-1 1.72e-1 1.24e-1 3.09e-1
PubMed 7.30e-3 6.51e-3 2.97e-2 1.69 1.55e-1 8.34e-3 2.73e-2 8.72e-2
C.2 Extended Persistence Image
Table 6: Mean squared error (MSE) for prediction of extended persistence image with vertex degree as node features (lower is better). The first group of three ER graphs are created with $n=100$ nodes and the second is created using $n=200$ nodes.
Graph Extended Persistence Image
DYMAG PDGNN GNN GAT GIN GRAND GRAND++
(Heat) (Wave) (Sprott)
$\text{ER}(p=0.04)$ 1.48e-2 $\pm$ 7.0e-4 6.37e-3 $\pm$ 4.2e-5 6.63e-1 $\pm$ 1.8e-3 3.07e-3 $\pm$ 1.3e-5 7.83e-1 $\pm$ 1.3e-2 3.82 $\pm$ 4.6e-2 4.83e-1 $\pm$ 7.1e-3 7.72e-1 $\pm$ 6.4e-3 2.39 $\pm$ 1.3e-1
$\text{ER}(p=0.06)$ 8.65e-3 $\pm$ 2.0e-4 2.79e-3 $\pm$ 1.6e-4 6.24e-1 $\pm$ 2.3e-3 1.37e-3 $\pm$ 2.4e-4 7.35e-1 $\pm$ 6.2e-3 1.57 $\pm$ 4.4e-2 4.09e-1 $\pm$ 5.2e-3 5.72e-1 $\pm$ 8.0e-3 1.15 $\pm$ 4.2e-2
$\text{ER}(p=0.08)$ 8.82e-3 $\pm$ 1.6e-3 2.54e-3 $\pm$ 6.1e-5 5.71e-1 $\pm$ 1.6 e-3 1.90e-3 $\pm$ 1.0e-4 9.46e-1 $\pm$ 4.0e-3 6.65e-1 $\pm$ 9.3e-3 3.92e-2 $\pm$ 5.6e-3 1.07e-2 $\pm$ 3.1e-3 8.59e-1 $\pm$ 7.4e-2
$\text{ER}(p=0.04)$ 8.91e-3 $\pm$ 1.2e-3 5.18e-3 $\pm$ 7.9e-5 8.04e-2 $\pm$ 7.9e-4 3.24e-3 $\pm$ 3.1e-5 6.73e-1 $\pm$ 1.4e-3 1.93e-1 $\pm$ 6.3e-3 3.72e-2 $\pm$ 1.0e-3 7.31e-2 $\pm$ 1.6e-3 2.85e-1 $\pm$ 4.4e-3
$\text{ER}(p=0.06)$ 7.41e-3 $\pm$ 9.2e-5 4.76e-3 $\pm$ 5.6e-5 1.39e-1 $\pm$ 2.3e-3 4.30e-3 $\pm$ 1.8e-4 7.29e-1 $\pm$ 1.3e-3 5.48e-1 $\pm$ 2.2e-3 3.69e-2 $\pm$ 5.0e-4 8.39e-2 $\pm$ 2.6e-3 3.28e-1 $\pm$ 8.9e-3
$\text{ER}(p=0.08)$ 4.57e-3 $\pm$ 1.7e-4 1.62e-3 $\pm$ 4.1e-5 1.35e-1 $\pm$ 1.7e-3 1.12e-3 $\pm$ 3.4e-5 1.28 $\pm$ 9.4e-2 1.87e-1 $\pm$ 5.5e-3 3.96e-2 $\pm$ 1.0e-3 4.12e-2 $\pm$ 7.2e-4 3.48e-1 $\pm$ 1.7e-3
Cora 1.45e-4 $\pm$ 3.8e-6 7.41e-5 $\pm$ 4.2e-7 3.51e-3 $\pm$ 4.1e-5 5.80e-5 $\pm$ 3.3e-7 2.72e-3 $\pm$ 7.2e-5 5.04e-2 $\pm$ 3.9e-4 5.53e-3 $\pm$ 1.8e-5 8.13e-4 $\pm$ 1.4e-5 8.16e-3 $\pm$ 3.7e-5
Citeseer 3.94e-4 $\pm$ 6.2e-6 1.58e-4 $\pm$ 7.1e-6 7.64e-3 $\pm$ 5.7e-5 1.29e-4 $\pm$ 3.0e-6 6.83e-3 $\pm$ 5.2e-5 2.93e-2 $\pm$ 1.0e-1 3.94e-3 $\pm$ 6.4e-5 6.87e-4 $\pm$ 9.2e-6 4.21e-3 $\pm$ 3.8e-5
PubMed 7.93e-3 $\pm$ 3.6e-5 3.29e-3 $\pm$ 8.8e-5 8.96e-2 $\pm$ 3.8e-4 2.74e-3 $\pm$ 1.7e-5 4.29e-1 $\pm$ 2.7e-3 4.79e-1 $\pm$ 3.0e-3 1.27e-1 $\pm$ 4.5e-3 3.37e-2 $\pm$ 5.1e-4 3.07e-1 $\pm$ 2.8e-3
To predict the extended persistence image, an MLP was applied to graph level representations (Equation 10). For baseline methods, the models were given the node degree as input features.
We present the results in the main body of the text using node degree as the filtration function in Table 6 based on 3 replicates. We experiment with the use of a different filtration function to generate the extended persistence image. The following results were obtained by using the clustering coefficient of nodes instead of node degree as the filtration parameter. The cluster coefficient is defined as:
$c(v)=\frac{2T(v)}{\text{deg}(v)(\text{deg}(v)-1)}$
where $T(v)$ is the number of triangles through node $v$ and $\text{deg}(v)$ is the degree of node $v$. Compared to the previous result, the higher MSE values here indicate that the topological properties of this filtration are harder to predict.
Table 7: MSE for prediction of extended persistence image with clustering coefficient as the filtration function. Vertex degree is utilized as the input node feature.
Dataset Extended Persistence Image
DYMAG PDGNN GNN GAT GIN GRAND GRAND++
(Heat) (Wave)
Cora 2.48 13.3 9.38 4.20 9.25 9.89 13.1 15.1
Citeseer 7.35 2.45 1.94 12.6 4.45 7.34 10.8 6.49
PubMed 1.28 6.04 4.52 7.94 7.50 2.17 6.07 4.75